A soild body starts rotating about stationary axis with an angular acceleration=t/50,t is in seconds.How soon after the beginning of rotation will the total acceleration vector of an arbitary point of the body form an angle 60degrees with its velocity vector?

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  krishna ,

 a  =  t/50

   = dw/ dt

hence

dw =  (t/50 )dt

on integrating

  w =   t2 /100


a tangential = ar = 2*10-2 t

a centripetal =w 2r = 10-4 t4

also

tan 60 =  √3 = a centripetal / a tangential

                  = t3 /  (2 * 102 )

hence

t3  =  2* 3*102
     346.4

or

t     =    7 seconds 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD